# Conformal compactification and affine-null metric formulation of the Einstein equations

### v1

@@ I Introduction
The concept of conformal compactification by Penrose [1, 2] has led to a thorough understanding of the structure and radiative properties of the spacetime of General Relativity. It gives mathematical meaning to paradoxical phrases like ‘points at infinity’, ‘close to infinity’, ‘a coordinate patch at infinity’ or alike. The meaning is achieved by adding a boundary $\mathcal{I}$ to the spacetime manifold $\tilde{\mathcal{M}}$, which contains the start and endpoints of all null geodesics that are in $\tilde{\mathcal{M}}$. The union of the spacetime manifold and this boundary surface gives rise to a bigger unphysical manifold $\mathcal{M}$ and it is accomplished by introducing a sufficiently differentiable non-negative scalar function $\Omega$ in the unphysical manifold while conformally relating the metric of the unphysical spacetime $g_{ab}$ with the physical metric $\tilde{g}_{ab}$, like $g_{ab}=\Omega^{2}\tilde{g}_{ab}.$ The triple $(\mathcal{M},g_{ab},\Omega)$ is the conformally compactified spacetime of the physical spacetime (~$(\tilde{\mathcal{M}},\tilde{g}_{ab})$. From the conformal metric relation, it is seen immediately that null vectors in the physical spacetime are also null vectors in the unphysical spacetime. Regarding the previous paradoxical statements, ‘points at infinity’ are now points on the boundary, while points ‘near infinity’ are in a half-open patch of the unphysical conformal spacetime, which includes the boundary and points not on the boundary. In addition, sufficient differentiability assumptions need to be made in order to evaluate geometric objects like connection and curvature and to assure a connectivity between boundary and the physical spacetime. The mathematical aspects of this ‘linking’ process between the conformal boundary and physical spacetime was studied in great detail by numerous authors, see e.g. review by [3]. Indeed, the relationship between physical quantities at finite distances in vacuum spacetimes and their connection to invariant quantities defined in conformal spacetimes was studied for applications in numerical relativity by [4, 5, 6, 7].
@@
Remaining on the classical side of studies, conformal methods play a role in the understanding of the nature of the singularities raising from the solution of Einstein equations [19, 20]. Moreover, the attachment of the boundary to the physical spacetime allows a bonafide investigation of the regularity of Einstein’s equations and their solutions using the conformal Einstein equations [21, 22, 23, 24, 25]. Based on those works, numerical relativists [26, 27, 28, 29] used the conformal Einstein equations to find numerical solutions to these highly non-linear partial differential equations; the articles [30, 31, 32] collect contibutions. These numerical methods often use the hyperboloidal slicing [33, 34, 35], which are spacelike hypersurfaces that become null in the limit towards the conformal boundary [36]. A recent software package to solve the conformal Einstein equations employing state-of-the-art numerical methods is the COFFEE package [37] and it was applied to find numerical solutions at both future and past null infinity [38].
@@
Numerical schemes in the Bondi-Sachs formalism base their foundations in the analytical considerations of [9, also see [42]] for vaccum spacetimes, who have shown that there is a patch containing null infinity in a conformal spacetime, in which exists a Bondi mass loss formula. Proper first numerical investigations in a Bondi-Sachs formalism were done in the early 1980ies by [43], who compactified the area distance coordinate of the Bondi metric to map null infinity on the numerical grid. This coordinate compactification is now the ‘standard way’ to reach null infinity in characteristic codes employing versions of the Bondi-Sachs metric. Various milestones and achievements have been reached with numerical characteristic Bondi-Sachs codes, for example, Cauchy Characteristic extraction of gravitational waves [44], the first long term stable evolution of a perturbed black hole [45], core collapse of relativisitic stars [46, 47], demonstration of the Choptuik critical phenomena [48, 49, 50] at null infinity [51], and Cauchy-Characteristic matching [52]; for reviews consult [53, 54] or for a listing of most recent works [52].
@@
An alternative way to the classical Bondi-Sachs metric (or versions thereof) for developing a metric based numerical algorithm is to use an affine-null metric [55]. The difference to the Bondi-Sachs metric is that the coordinate along the null generators is an affine parameter instead of an area distance. The affine parameter coordinate has the advantage that it remains regular at an apparent horizon, where the expansion rate of outgoing null rays vanishes and the area distance coordinate renders the volume element of its defining metric infinite. Using an affine parameter coordinate in a null metric is beneficial when simulating a forming black hole or in situations of near horizon formation, where the area distance leads to numerical inaccuracies due to its singular behavior [56]. However, the prize to pay with an affine-null metric is that the nice and convenient Bondi-Sachs hierarchy of its resulting partial differential equations is broken due to the presence of an additional time derivative in one of the equations. Nevertheless, a hierarchical system of hypersurface-evolution equations can be restored via some rather non-trivial variable transformations [55, 57, 58, 59]. Recent numerical investigations [60] in physical space, also employed a compactified version of the affine parameter to reach null infinity and also extracted the gravitational redshift and Bondi mass decay of supercritical solutions forming a black hole. In this work we show how these extracted regularized quantities (including the scalar field) can naturally be obtained from well defined fields at null infinity in the sense of a conformal compactification. We calculate the conformal field equations in affine-null coordinate in a conformally compactified spherically symmetric spacetime filled with a massless scalar field. More importantly, we will demonstrate how to successfully obtain a hierarchical set of equations in the conformal metric, utilizing field variables that are regular and well-defined throughout the entire non-physical spacetime, including its future null boundary. The obtained conformal field equations are then compared with those resulting from a coordinate transformation in physical space. During the course of this work we also derive solution of the affine-null metric equations in the neighborhood the conformal boundary $\mathcal{I}$, which we have not seen in this context.
@@
In Sec. II, we present the conformal field equations in hierachical form and give their solution at $\mathcal{I}$. For comparison, in Sec. III, we present the hypersurface-evolution equations with a compactified physical affine parameter, while no conformal boundary is introduced. Our results are discussed with a numerical example in Sec. IV. In App. A, we repeat for completeness the analysis of Sec. II and III using the classical Bondi-Sachs metric. In App. B, we derive a general representation of the Schwarzschild metric in conformal affine-null formulation serving for comparison with its Bondi (inertial) counterpart.
@@
We use units with $c=1$ and label the gravitational constant with $\kappa$. The metric signature is $+2$ and we denote with a tilde coordinates and tensors (like metric, scalar field, Ricci tensor, covariant derivative, etc) in physical space, while those quantities without a tilde are in conformal space. This nomenclature is the one of [2] and [3].
@@ II Conformal compactification and field equations
We consider a four dimensional smooth physical spacetime (~$(\tilde{\mathcal{M}},\tilde{g}_{ab})$ filled with a massless physical scalar field $\tilde{\Phi}$. The spacetime is charted with associated coordinates $\tilde{x}^{a}$ and the covariant derivative with respect to $\tilde{g}_{ab}$ is ~$\tilde{\nabla}_{a}$. The field equations are the Einstein scalar field equations
@@
To relate the physical spacetime (~$(\tilde{\mathcal{M}},\tilde{g}_{ab})$ and the unphysical spacetime (~$(\tilde{\mathcal{M}},\tilde{g}_{ab})$, the cooresponding unphysical metric $g_{ab}$ is defined via the conformal relation
@@
Since $\Omega=0$ and $\nabla_{a}\Omega\neq 0$ on $\mathcal{I}$, a finite value of $\Phi$ on $\mathcal{I}$ requires that $\tilde{\Phi}$ is sufficiently differentiable in a coordinate patch near $\mathcal{I}$ and that $\tilde{\Phi}$ vanishes at in the limit towards $\mathcal{I}$. In other words, the physical field $\tilde{\Phi}$ requires a local smooth extension from $\mathcal{M}/\mathcal{I}$ to $\mathcal{I}$ [3]. Unphysical coordinates in $\mathcal{M}$ are denoted with $x^{a}$ and the covariant derivative with respect to $g_{ab}$ is $\nabla_{a}$.
@@
Using the field equations (1) in physical space we find the field equations for the unphysical spacetime $(\mathcal{M},g_{ab})$
@@
determined by an area distance $\mathcal{R}$ in the unphysical conformal space time $(\mathcal{M},\Omega,g_{ab})$.
@@ II.1 Conformal compactification with affine-null coordinates in spherical symmetry
According to the twice contracted Bianchi identies and the local conservation of energy-momentum the first equation, (LABEL:eq:supp_conf_an), is the supplementary equation. In a numerical algorithm, it is used to fix the boundary conditions of the fields. The equations (20b)-(20d) are the main equations used to evolve the fields. But due to the presence of the $\mathcal{R}_{,u}$ derivative, equations (20b)-(20d) are not suited for numerical work, since they do have not the typical hierarchy due to the choice of a characteristic foliation like e.g. in the Bondi-Sachs case [40, 41, 9, 54].
@@ II.2 Regularity condition of the fields at the vertex of the null cone and the respective field equations
If the system (28) is solved for $x\geq 0$ and $\mathcal{R}\geq 0$, it has still singularities at the conformal areal distance $\mathcal{R}=0$. Then $\mathcal{R}=0$ is a vertex of a null cone and the coordinate singularity is due to the singular nature of the vertex. It can then be dealt with the requirements of the Fermi normal coordinate system (imposed in the physical spacetime, see e.g. [67, 65]) at $x=0$, which imply
@@ III Coordinate compactification in physical space
where $\tilde{\lambda}(0)=$ and $\lim_{\hat{x}\rightarrow 1}\tilde{\lambda}=\infty$. This means the physical domain of $\tilde{\lambda}$ is completely mapped onto the domain of $\hat{y}$ and is referred to as coordinate domain of $\tilde{\lambda}$ is compactified.
@@
IV Summary and discussion

### v2

@@ I Introduction
The concept of conformal compactification by Penrose [1, 2] has led to a thorough understanding of the structure and radiative properties of the spacetime of General Relativity. It gives mathematical meaning to paradoxical phrases like ‘points at infinity’, ‘close to infinity’, ‘a coordinate patch at infinity’ or alike. The meaning is achieved by adding a boundary $\mathcal{I}$ to the spacetime manifold $\tilde{\mathcal{M}}$, which contains the start and endpoints of all null geodesics that are in $\tilde{\mathcal{M}}$. The union of the spacetime manifold and this boundary surface gives rise to a bigger unphysical manifold $\mathcal{M}$ and it is accomplished by introducing a sufficiently differentiable non-negative scalar function $\Omega$ in the unphysical manifold, while conformally relating the metric of the unphysical spacetime $g_{ab}$ with the physical metric $\tilde{g}_{ab}$, like $g_{ab}=\Omega^{2}\tilde{g}_{ab}.$ The triple $(\mathcal{M},g_{ab},\Omega)$ is the conformally compactified spacetime of the physical spacetime (~$(\tilde{\mathcal{M}},\tilde{g}_{ab})$. From the conformal metric relation, it is seen immediately that null vectors in the physical spacetime are also null vectors in the unphysical spacetime. Regarding the previous paradoxical statements, ‘points at infinity’ are now points on the boundary, while points ‘near infinity’ are in a half-open patch of the unphysical conformal spacetime, which includes the boundary and points not on the boundary. In addition, sufficient differentiability assumptions need to be made in order to evaluate geometric objects like connection and curvature and to ensure a connectivity between boundary and the physical spacetime. The mathematical aspects of this ‘linking’ process between the conformal boundary and physical spacetime was studied in great detail by numerous authors, see e.g. review by [3]. Indeed, the relationship between physical quantities at finite distances in vacuum spacetimes and their connection to invariant quantities defined in conformal spacetimes was studied for applications in numerical relativity by [4, 5, 6, 7].
@@
Remaining on the classical side of studies, conformal methods play a role in the understanding of the nature of the singularities arising from the solution of Einstein equations [19, 20]. Moreover, the attachment of the boundary to the physical spacetime allows a bonafide investigation of the regularity of Einstein’s equations and their solutions using the conformal Einstein equations [21, 22, 23, 24, 25]. Based on those works, numerical relativists [26, 27, 28, 29] used the conformal Einstein equations to find numerical solutions to these highly non-linear partial differential equations; the articles [30, 31, 32] collect contributions. These numerical methods often use the hyperboloidal slicing [33, 34, 35], which are spacelike hypersurfaces that become null in the limit towards the conformal boundary [36]. A recent software package to solve the conformal Einstein equations employing state-of-the-art numerical methods is the COFFEE package [37] and it was applied to find numerical solutions at both future and past null infinity [38].
@@
Numerical schemes in the Bondi-Sachs formalism base their foundations in the analytical considerations of [9, also see [42]] for vaccum spacetimes, who have shown that there is a patch containing null infinity in a conformal spacetime, in which exists a Bondi mass loss formula. The first proper numerical investigations in a Bondi-Sachs formalism were done in the early 1980ies by [43], who compactified the area distance coordinate of the Bondi metric to map null infinity on the numerical grid. This coordinate compactification is now the ‘standard way’ to reach null infinity in characteristic codes employing versions of the Bondi-Sachs metric. Various milestones and achievements have been reached with numerical characteristic Bondi-Sachs codes, for example, Cauchy Characteristic extraction of gravitational waves [44], the first long term stable evolution of a perturbed black hole [45], core collapse of relativisitic stars [46, 47], demonstration of the Choptuik critical phenomena [48, 49, 50] at null infinity [51], and Cauchy-Characteristic matching [52]; for reviews consult [53, 54] or for a listing of most recent works [52].
@@
An alternative way to the classical Bondi-Sachs metric (or versions thereof) for developing a metric based numerical algorithm is to use an affine-null metric [55]. The difference to the Bondi-Sachs metric is that the coordinate along the null generators is an affine parameter instead of an area distance. The affine parameter coordinate has the advantage that it remains regular at an apparent horizon, where the expansion rate of outgoing null rays vanishes and the area distance coordinate renders the volume element of its defining metric infinite. Using an affine parameter coordinate in a null metric is beneficial when simulating a forming black hole or in situations of near horizon formation, where the area distance leads to numerical inaccuracies due to its singular behavior [56]. However, the price to pay with an affine-null metric is that the nice and convenient Bondi-Sachs hierarchy of its resulting partial differential equations is broken due to the presence of an additional time derivative in one of the equations. Nevertheless, a hierarchical system of hypersurface-evolution equations can be restored via some rather non-trivial variable transformations [55, 57, 58, 59]. Recent numerical investigations [60] in physical space, also employed a compactified version of the affine parameter to reach null infinity and also extracted the gravitational redshift and Bondi mass decay of supercritical solutions forming a black hole. In this work we show how these extracted regularized quantities (including the scalar field) can naturally be obtained from well defined fields at null infinity in the sense of a conformal compactification. We calculate the conformal field equations in affine-null coordinate in a conformally compactified spherically symmetric spacetime filled with a massless scalar field. More importantly, we will demonstrate how to successfully obtain a hierarchical set of equations in the conformal metric, utilizing field variables that are regular and well-defined throughout the entire non-physical spacetime, including its future null boundary. The obtained conformal field equations are then compared with those resulting from a coordinate transformation in physical space. During the course of this work we also derive a solution of the affine-null metric equations in the neighborhood of the conformal boundary $\mathcal{I}$, which we have not seen in this context.
@@
In Sec. II, we present the conformal field equations in hierachical form and give their solution at $\mathcal{I}$. For comparison, in Sec. III, we present the hypersurface-evolution equations with a compactified physical affine parameter, while no conformal boundary is introduced. In Sec. IV, we apply our results to extract the scalar news function at null infinity. Our results are discussed in Sec. V. In App. A, we repeat for completeness the analysis of Sec. II and III using the classical Bondi-Sachs metric. In App. B, we derive a general representation of the Schwarzschild metric in conformal affine-null formulation serving for comparison with its Bondi (inertial) counterpart.
@@
We use units with $G=1$, $c=1$ and label the gravitational constant with $\kappa$. The metric signature is $+2$ and we denote with a tilde coordinates and tensors (like metric, scalar field, Ricci tensor, covariant derivative, etc) in physical space, while those quantities without a tilde are in conformal space. This nomenclature is the one of [2] and [3].
@@ II Conformal compactification and field equations
We consider a four dimensional smooth physical spacetime (~$(\tilde{\mathcal{M}},\tilde{g}_{ab})$ filled with a massless physical scalar field $\tilde{\Phi}$. The spacetime is charted with associated coordinates $\tilde{x}^{a}$ and the Levi-Civita covariant derivative with respect to $\tilde{g}_{ab}$ is ~$\tilde{\nabla}_{a}$. The field equations are the Einstein scalar field equations
@@
To relate the physical spacetime (~$(\tilde{\mathcal{M}},\tilde{g}_{ab})$ and the unphysical spacetime, the corresponding unphysical metric $g_{ab}$ is defined via the conformal relation
@@
Since $\Omega=0$ and $\nabla_{a}\Omega\neq 0$ on $\mathcal{I}$, a finite value of $\Phi$ on $\mathcal{I}$ requires that $\tilde{\Phi}$ is sufficiently differentiable in a coordinate patch near $\mathcal{I}$ and that $\tilde{\Phi}$ vanishes at in the limit towards $\mathcal{I}$. In other words, the physical field $\tilde{\Phi}$ requires a local smooth extension from $\mathcal{M}/\mathcal{I}$ to $\mathcal{I}$ [3]. Unphysical coordinates in $\mathcal{M}$ are denoted with $x^{a}$ and the Levi-Civita covariant derivative with respect to $g_{ab}$ is $\nabla_{a}$.
@@
Using the field equations (1) in physical space we find the field equations for the unphysical spacetime $(\mathcal{M},g_{ab},\Omega)$
@@
determined by an area distance $\mathcal{R}$ in the unphysical conformal spacetime $(\mathcal{M},g_{ab},\Omega)$.
@@ II.1 Conformal compactification with affine-null coordinates in spherical symmetry
According to the twice contracted Bianchi identities and the local conservation of energy-momentum the first equation, (LABEL:eq:supp_conf_an), is the supplementary equation. In a numerical algorithm, it is used to fix the boundary conditions of the fields. The equations (20b)-(20d) are the main equations used to evolve the fields. But due to the presence of the $\mathcal{R}_{,u}$ derivative, equations (20b)-(20d) are not suited for numerical work, since they do have not the typical hierarchy due to the choice of a characteristic foliation like e.g. in the Bondi-Sachs case [40, 41, 9, 54].
@@ II.2 Regularity condition of the fields at the vertex of the null cone and the respective field equations
If the system (28) is solved for $x\geq 0$ and $\mathcal{R}\geq 0$, it still has singularities at the conformal areal distance $\mathcal{R}=0$. Then $\mathcal{R}=0$ is a vertex of a null cone and the coordinate singularity is due to the singular nature of the vertex. It can then be dealt with the requirements of the Fermi normal coordinate system (imposed in the physical spacetime, see e.g. [67, 65]) at $x=0$, which imply
@@ III Coordinate compactification in physical space
where $\tilde{\lambda}(0)=$ and $\lim_{\hat{x}\rightarrow 1}\tilde{\lambda}=\infty$. This means the physical domain of $\tilde{\lambda}$ is completely mapped onto the domain of $\hat{x}$ and is referred to as coordinate domain of $\tilde{\lambda}$ is compactified.
@@
IV Direct extraction of scalar News function at null infinity
In the study of gravitational collapse and scalar field dynamics near criticality, an important quantity is the scalar news function, which governs the radiative loss of mass through null infinity. Accurate extraction of this quantity is essential for tracking the decay of the Bondi mass and understanding the global properties of the spacetime. As an application of our previous findings about the partial differential equations in physical and conformal spacetimes, we revisit a key technical aspect of previous work [60].
Therein, the news calculation on the compactified grid requires evaluating a second-order mixed $u\hat{x}$-derivative of the scalar field. This numerical derivative was noisy in [60], due to the employed spectral method and the e-folding [49] of the scalar field near the critical solution. For this reason, the news function driving the Bondi mass decay was not extracted. However, the use of the regularized scalar field $\Phi$ (see (64)) avoids the calculation of this mixed derivative. In fact the news function can be read off directly from the hypersurface fields given by the right hand side of (65d) at the outer boundary.
Taking these considerations into account, we now demonstrate how the numerical code of [60] can be adapted to compute the news function. The code solves the field equations with the origin, $\tilde{\lambda}=0$, on to the central worldline of spherical symmetry, which implies that the metric boundary conditions satisfy the regularity requirements $r=\tilde{\lambda}+O(\tilde{\lambda}^{3})$ and $V=1+O(\tilde{\lambda}^{2})$. As a consequence, the second-order differential equation of the $V$-hypersurface equation can be cast into a first-order one for the new auxiliary field $Q=(V-1)/\tilde{\lambda}$. A further consequence is $Z_{0}=0$. We evolve the initial data ${\tilde{\Phi}=\sqrt{12}\left[\arcsin(\frac{b\tilde{\lambda}}{a+\tilde{\lambda}})-\arcsin b\right]}$ with $a=2$ and $b\simeq 0.9472018$. The given parameter choice corresponds to supercritical initial data evolving into a black hole spacetime for code time [^1] For the employed initial data, the critical solution is found for $b=b_{c}\simeq 0.947201675$. The parameter $a$ is a scaling parameter. It determines the time, $u_{c}$, of formation of the critical solution as time $u_{c}=3a$. Black holes thus form for values $b>b_{c}$ and at times bigger than $u_{c}$, see [60]. $\tilde{u}\gtrsim 6$.
The physical affine parameter is substituted by the grid function [^2] This is the compactified grid function of [60] with the parameter choice $A=1$. $\tilde{\lambda}=2(1+\tilde{x})(1-\tilde{x})^{-1}$ with $\tilde{x}\in[-1,1]$ so that $\tilde{\lambda}(-1)=0$ and $\lim_{\tilde{x}\rightarrow 1}\tilde{\lambda}(\tilde{x})=\infty.$ The code of [60] uses the physical scalar field $\tilde{\Phi}$ and already employs the renormalized area distance $R=(1-\tilde{x})r$. We change the implementation of [60] by using ${\Phi=(1-\tilde{x})\tilde{\Phi}}$ instead of $\tilde{\Phi}$. The physical relevant fields, $f\in(\Phi,R,L,Q,V)$, are discretized like $f(u_{n},\tilde{x}_{i})=\sum_{k=0}^{N}f_{k}(u_{n})T_{k}(\tilde{x}_{i})$, where $N+1$ is the number of grid points discretizing equidistantly the half unit circle in $N$ arc segments, $u_{n}=n\Delta u$ with $\Delta u=(N+1)^{-2}$ is the discretized representation of the proper time along the central geodesic, $\tilde{x}_{i}$ are Gauss-Lobatto points, ${\tilde{x}_{i}=-\cos(\frac{i\pi}{N})}$ with $i=0,...N$, which are the extrema of the Chebycheff polynomials $T_{k}(\tilde{x})$ of first kind. Under these definitions the modified system gives the set of hypersurface equations
$\tilde{\Phi}_{,\tilde{x}}$ $=$ $(1-\tilde{x})\Phi_{,\tilde{x}}-\Phi$ (66a)
00 0 $=$ $R_{,\tilde{x}\tilde{x}}+\frac{1}{2}\kappa R\Big{(}\tilde{\Phi}_{,\tilde{x}}\Big{)}^{2},$ (66b)
${L}_{,\tilde{x}}$ $=$ $\frac{2(1+\tilde{x})\tilde{\Phi}_{,\hat{x}}}{R},$ (66c)
$Q_{,\tilde{x}}$ $=$ $\frac{1}{(1+\tilde{x})^{2}}-\frac{4}{R^{2}}+\frac{1}{2}\left(\frac{L}{1+\tilde{x}}\right)^{2},$ (66d)
and the evolution equation
$\Phi_{,u}=\frac{1}{2}\left\{\frac{L}{R}+\frac{1}{4}[(1-\tilde{x})+2(1+\tilde{x})Q]\tilde{\Phi}_{,x}\right\}.$ (67)
The system is solved by presenting initial values for $\Phi$ on an initial null cone $u=0$, while using the boundary conditions $R(u,-1)=0$, $R_{,\tilde{x}}(u,-1)=2$ and $L(u,-1)=Q(u,-1)=0$, while $\Phi_{,u}(u,-1)=\Phi_{,\tilde{x}}(u,-1)$. With the initial data on the cone $u=0$, the hypersurface equations (66) are solved sequentially. Then the right hand side of (67) can be solved on $u=0$. Given the time step $\Delta u=(N+1)^{-2}$, the scalar field is then calculated with the Shu and Osher third order stability [69] preserving Runge-Kutta scheme to find $\Phi(\Delta u)$. The scalar field $\Phi(\Delta u)$ on the cone $\Delta u$ then serves as new initial data to obtain $\Phi(2\Delta u)$. This hierarchical hypersurface-evolution algorithm is repeated until the end of the simulation time. In Fig. 1, we show the convergence of spatial distribution of the absolute error (with respect to a high resolution run) of the scalar field. If $\Delta u$ is divided by a factor of two [^3] This corresponds to the number of time steps minus one being doubled. while keeping the final simulation time fixed, the error decays by a factor of $8=2^{3}$. This is consistent with third order convergence. The inlet in Fig. 1 also confirms that the error of the time integration is $O(\Delta u^{3})$.
Figure 1: Absolute error of the scalar field $\Phi$ as a function of $\tilde{x}$ resulting from six different temporal resolution runs with $N_{u}\in\{{6.3\cdot 10^{3}},{1.3\cdot 10^{4}},{2.5\cdot 10^{4}},{5\cdot 10^{4}},1\cdot 10^{5},2\cdot 10^{5}\}$ time steps. All errors are calculated at the same time $u=6.13$. The curve with the blue stars, orange squares, green diamonds, red crosses and violet dots correspond absolute error difference with respect to the run having $2\cdot 10^{5}$ time steps with the run with $6.3\cdot 10^{3}$, ${1.3\cdot 10^{4}}$, ${2.5\cdot 10^{4}}$, ${5\cdot 10^{4}}$ and $1\cdot 10^{5}$ time steps, respectively. The spatial number of grid points is kept fixed at $33$ collocation points. By doubling the number of temporal spacings the error decays by a factor of $8=2^{3}$. The inlet curve with blue squares shows the dependence of the $\ell_{2}$-norm at final time as function of the time step employed in the simulations. The orange dotted curve displays the expected behavior of the temporal discretization error $\sim\Delta u^{3}$. The alignment of the $\ell_{2}$-norm values to this curve confirms the third order convergence of the numerical scheme.
For the massless Einstein-scalar field equations in spherical symmetry, the variables of physical interest at null infinity are: i) the redshift $H$, ii) the Bondi time $u_{b}$, iii) the scalar monopole $C$, iv) the Bondi mass $m_{b}$, and v) the news function $N$. The function $H$ is numerically extracted at the outer boundary according to $H(u)=R(u,-1)/4$. Given $H$, the Bondi time is found by integration of (50). Moreover, once $H$ is determined, the Bondi mass is found from $m_{B}(u)=H(u)-R_{,x}(u,1)/2$. Given $H$, the scalar monopole $C$ is found numerically by evaluation of the scalar field at $\tilde{x}=1$, i.e. $C=\Phi(u,-1)/(4H)$. The news function $N$ is defined with respect to the Bondi time $\tilde{u}_{b}$ in (52). In terms of the code time $u$, it is $N=HC_{,u}=HH_{,u}\Phi_{[0]}+H^{2}\Phi_{[0],u}$ which follows from (47), (50) and (52). The derivative $\Phi_{[0],u}$ can be read from the right hand side of (67) evaluated at the outer boundary, while the derivative $H_{,u}$ is read from the right hand side of (66d). The latter is a consequence of $\lim_{\lambda\rightarrow\infty}Q=(\ln H^{2})_{,u}$, see [58]. With knowledge of the news function, Bondi mass and Bondi time, the asymptotic balance law (51) can be calculated to verify the consistency of the extracted quantities.
Figure 2: In all three panels, quantities are evaluated at null infinity, $\tilde{x}=1$, and the abscissa is the Bondi time $u_{b}$. The upper panel A shows the evolution of the redshift factor $H$ (red dashed line) and the Bondi mass $m_{B}$ (solid black line). The middle panel B shows the news function $N$ (dashed red line) and the scalar monopole $C$ (solid black line). The lower panel C shows the numerical error of the calculation of the right hand side of (51). The vertical dotted lines correspond to the Bondi time of the last null cone connecting with the central geodesic. The horizontal dotted line in panel A and the blue star point out the final value of the Bondi mass at the threshold of black hole formation.
In Fig. 2, we display the behavior of the asymptotically relevant variables, redshift $H$, Bondi mass $m_{B}$, scalar monopole $C$, news function $N$ and the balance law (51) as a function of the Bondi time $u_{b}$. To our knowledge, this is the first extraction of a news function in the context of critical phenomena while using affine-null coordinates. Up until now scalar field news functions have been extracted using Bondi-Sachs coordinates by [51, 70, 71]. As explained in Sec. II, the formation of a black hole is indicated if the redshift $H\rightarrow 0$. In panel A of Fig. 2, it is seen that the redshift (red dashed line) drops to zero at the Bondi time $u_{H}:=u_{b}\sim 26.70$. This $\tilde{u}_{b}=const$ slice corresponds to the last null cone that connects an asymptotic observer at null infinity with the central geodesic, where $\tilde{x}=-1$ and from where the system (66) and (67) is integrated towards $\tilde{x}=1$. Beyond the time $u_{H}$, an asymptotic observer has no information on how the scalar field continues to collapse toward the central geodesic.
Figure 3: Evolution of the scalar field and its $u-$derivative at the origin in the compactified coordinate where $\tilde{\lambda}=0$. The dotted line corresponds to the code time of the last null cone connecting to $\mathcal{I}$.
Panel A also displays the (exponential) decay of the Bondi mass with a final value (blue star) of $m_{B}=1.97\cdot 10^{-2}$ measured at $\mathcal{I}$. The news function (red dashed line) and the scalar monopole (black solid line) are shown in panel B of Fig. 2. In the behavior of the two functions one can see that the news function is the derivative of the monopole with respect to the Bondi time $u_{b}$. It can be understood by correlating the zeros of the news function with the extrema of the scalar monopole. We observe in panel B that they coincide. The Bondi mass decay is coupled with the square of the news function via (51). In panel B of Fig. 2 the right hand side of (51) is shown using the data of the Bondi mass displayed in panel A and the news function in panel B. This balance law was also shown graphically by displaying its components by Barreto [70] in physical space. We can see that the balance law holds up to numerical precision. For about 80% of the overall simulation time of $u_{H}$, the absolute error in the balance law is $\sim 10^{-7}$. It then increases to about $10^{-3}$ until the simulation crashes due to black hole formation. The increase of error can be understood by looking at the evolution of the scalar field at the origin, i.e along the central geodesic. Thereat are the start points of the null rays emanating to null infinity. The scalar field (black curve) and its $u$-derivative (red dashed curve) both evaluated at the origin are shown in Fig. 3. The abscissa is the proper time at the central geodesic and it is also the time measured in the numerical simulation. The central time determines the Bondi time in Fig. 2 by integrating (50). We observe that the zero crossings of the scalar field decrease exponentially towards the final time. The central time $u=6.13$ has the Bondi time $u_{H}$ at $\mathcal{I}$, which is the time when the last null cone emanating from the central is connected with null infinity. The exponential decrease in the zero crossings of the scalar field towards final time, also means that the $u$-derivative of the scalar field along the central geodesic exponentially in amplitude towards the final time (see the red curve in Fig. 3). As the $u$-derivative of the scalar field along the geodesic is determined by a $\tilde{\lambda}$-derivative evaluated at [^4] The system (59) and the regularity conditions for a freely falling observer at the central geodesic imply $\Phi_{,\tilde{u}}|_{\tilde{\lambda}=0}=\Phi_{,\tilde{\lambda}}|_{\tilde{\lambda}=0}$. Consequently, in code variables $\Phi_{,u}|_{\tilde{x}=-1}=\Phi_{,\tilde{x}}|_{\tilde{x}=-1}$, where the spatial derivative is taken over the data on a given cone $u=const$. $\tilde{\lambda}=0$ , for times close to black hole formation time, the spatial derivatives at the vertex also become stiffer with increasing amplitudes. As consequence high numerical resolution is required at the vertex to capture the scalar fields dynamical behavior along the geodesic. It is this stiffness and because of the employed spectral methods that leads to the increase of error towards $10^{-3}$ in panel C of Fig. 2 and spoils the numerical accuracy in the evaluation of the balance law at times before black hole formation. Despite the loss of accuracy, the temporal convergence (see Fig. 1) on the final null cone $u_{H}$ indicates the numerical consistency/convergence of the results.
V Summary and discussion